OpenStudy (ciarabelen):

Solve the following system of equations algebraically for ALL values of x and y. y=x^2+2x-8 y=2x+1

OpenStudy (zehanz):

Before you start, can you think of how many x-values possibly are solution of this system?

OpenStudy (ciarabelen):

3?

OpenStudy (anonymous):

there are 2 solutions to x, since it is squared.

OpenStudy (ciarabelen):

ok what do i do next?

OpenStudy (anonymous):

try and eliminate one variable. y is the easiest to eliminate since you have two equations for y

OpenStudy (ciarabelen):

ok now i have 0=x^2-9

OpenStudy (anonymous):

factor that.

OpenStudy (ciarabelen):

x= 3 or -3

OpenStudy (anonymous):

yes....y should be very easy to find now

OpenStudy (ciarabelen):

y=0?

OpenStudy (anonymous):

y = 2x +1 plug in the values of x that you found

OpenStudy (ciarabelen):

y=7

OpenStudy (anonymous):

there are 2 solutions to y since there are 2 solutions to x.

OpenStudy (ciarabelen):

y=7 and y=-7

OpenStudy (anonymous):

no.... y = 2(3) + 1 = 7 y = 2(-3) + 1 = -5

OpenStudy (ciarabelen):

oh ok thanks